### v1

# Supplementary Information

Abstract
Single-molecule spectroscopy (SMS) is an exceptionally sensitive technique, but its inherently limited photon budget produces noisy data that can readily lead to subjective analyses, fitting errors, and reduced statistical power, obscuring true subpopulations and their dynamics. Here, we present an unbiased, objective method to cluster two-dimensional single-molecule data and demonstrate it on fluorescence lifetime–intensity correlations. The clustering method is based on Gaussian mixture modeling, with the optimal number of clusters determined through the Bayesian information criterion (BIC). The BIC score per cluster, which displays in general a non-monotonically decreasing trend, presents multiple local minima as candidate solutions for the number of fitted clusters. We also demonstrate the usefulness of statistically grouping resolved levels. The clustering protocol was benchmarked on simulated data and applied to experimental data from the Alexa Fluor 647 dye, QD 605, and the main light-harvesting complexes of plants and cyanobacteria. The combined application of grouping and clustering achieves substantial noise reduction and the identification of relevant, physically meaningful states that would typically be obscured by manual inspection.
[…]
{sigstatement}
Single-molecule spectroscopy is uniquely capable of probing the structural dynamics, reaction kinetics, and environmental responses of individual molecules and nanoscaled objects, providing simultaneous insight into parameters such as fluorescence intensity, lifetime, and spectral behavior. However, the inherent sensitivity makes this technique highly susceptible to noise, which may obscure physically relevant subpopulations and their dynamics. Clustering is often employed to assist in identifying these subpopulations, but this is typically done on the basis of user input, which is subject to bias and limits reproducibility. Here, we present an unbiased clustering protocol that can operate on any 2D dataset, enabling consistent and objective identification of molecular subpopulations.
Introduction
The pioneering work of Moerner and Kador on single molecules (SMs) (moerner1989optical) paved the way for research into several highly sensitive optical imaging and spectroscopy techniques capable of detecting nanoscale molecular dynamics. Among these are SM Förster resonance energy transfer (smFRET) (roy2008practical, uphoff2010monitoring), time-resolved fluorescence methods that are usually based on time-correlated single photon counting (TCSPC) (gruber2018isolated, bockenhauer2011conformational), single-particle tracking techniques (van2022real, manzo2015review), superresolution localization imaging methods such as Photoactivated Localization Microscopy (PALM)  (hess2006ultra, betzig2006imaging) and Points Accumulation for Imaging in Nanoscale Topography (PAINT) (sharonov2006wide), SM electrophysiology techniques such as single-channel patch-clamp detection (neher1976single, harms2003probing) and nanopore electrophysiology (bayley2001stochastic), and SM force spectroscopy (yen2018improving). SM techniques are uniquely capable of detecting physically relevant subpopulations, which represent distinct functional states that the system can transition between, information that is masked or averaged out in ensemble techniques.
Raw SM experimental data can often be multidimensional depending on the type and purpose of the measurements. Each data point in a multidimensional parameter space could display information such as conformational dynamics, reaction kinetics, diffusion properties, electronic properties, and environmental responses, and these properties are often correlated with one another. The dimensionality of the data usually scales with computational complexity and visualization limitations. As such, subsets comprising only two or three parameters at a time are typically selected. Clustering can be effectively done on these lower-dimensional distributions and serves a dual purpose in denoising the data and isolating different subpopulations, some of which may be statistically rare but interesting. These data clusters may be physically linked to different functional states, for which the switching dynamics between these clusters can easily be determined. Examples include neural spike detection in patch-clamp detection (ghaderi2018electrophysiological), spectral clustering (kruger2017reduced), smFRET common trajectory identification (blanco2015single), and fluorescence intensity–lifetime correlation classification (schlau2015single).
The choice of clustering protocol depends on various factors, including the general observed trend, computational intensiveness, and prior knowledge. Data clustering is generally considered an unsupervised data classification method (although supervised and semi-supervised clustering also exist) whereby an algorithm identifies patterns in the dataset without prior ground-truth labels encoded into the algorithm. Even within this framework, a user can specify conditions under which the algorithm should operate—the maximum number of clusters, for instance—without altering its unsupervised nature. Although the number of clusters may be user-defined, the actual assignment of data points to clusters remains determined entirely by the algorithm, not by any user-supplied ground truth. Subjective clustering, as is the case for K-means and Gaussian clustering (schlau2015single, joshi2021single), may potentially introduce user bias, which could prevent reproducibility and negatively affect biological relevance and accuracy. Computationally intensive clustering, such as Bayesian cluster analysis (see, e.g., Refs. (rubin2015bayesian, griffie2016bayesian)), tends to be slow and may result in overfitting or underfitting the data. Although Refs. (schlau2015single, rubin2015bayesian, griffie2016bayesian) yielded excellent insight into the dynamics of the respective studied systems, the clustering algorithm used either calls for manual selection of the number of clusters, may be computationally expensive, or may overfit the data.
A more robust approach would involve employing a clustering algorithm that is scalable, probabilistic, and free from bias—a requirement rendered even more critical in light of the inherent complexities and limitations of data analysis. SM data involves a considerable amount of inherent noise that arises from sources such as autofluorescence of optical components, scattered light, other background signals, laser fluctuations, detector noise, sample inhomogeneities, and shot noise. This noise can easily shroud weak signals from rare events or transient states. Adding to these uncertainties are possibly non-optimized experimental conditions, which can cause additional experimental error.
Single-molecule spectroscopy (SMS) is an exquisitely sensitive way to detect temporal parameters such as fluorescence brightness, polarization, spectral properties (such as peak position or spectral linewidth), lifetime, and dwell times of a particular state and can be performed on virtually any nanoscale emitter. Examples of such emitters include quantum dots (QDs) (rombach2013blinking, palstra2021intermittency, palstra2021python), fluorescent dyes (zhang2021investigating), and dendrimer-based systems in which fluorophores can be embedded within their highly-branched polymer scaffolds (paulo2010single, kohn2001single). These systems generally exhibit binary switching behavior in their fluorescence intensity and lifetime, characterized by a well-defined bright, long-lifetime state, and a darker, quenched state, often called “on" and “off" states, respectively, a phenomenon known as photoblinking or fluorescence intermittency. However, many systems are known to deviate from such a bimodal behavior. Even single, fluorophore molecules have been shown to display multiple emissive states (wang2013lifetime). For QDs, the number of accessible states depends on their type, structure, and the excitation conditions. For example, type II-IV QDs generally display bimodal lifetime–intensity distributions (palstra2021python, efros2016origin), while perovskite QDs exhibit more continuous distributions (palstra2021intermittency, yarita2017impact). Clustering at the SM level can be a useful tool to decipher and interpret this complex fluorescence behavior (rombach2013blinking, munoz2022memory, pelton2007evidence, chu2024single).
The presence of multiple chromophores in multichromophoric systems contributes significantly to the heterogeneity in their data space. Examples of such systems include the aforementioned dendrimers as well as conjugated polymers, azobenzenes, diarylethenes, synthetic dye aggregates, DsRed, and photonic wires. Photosynthetic light-harvesting complexes form a unique group of multichromophoric systems because of the potential biological significance of data clusters. For example, a subtle change in protein conformation may switch many light-harvesting complexes between a light-harvesting and a photoprotective state (kruger2011conformational), and these transitions have been directly correlated with photoblinking (kruger2012controlled). In addition, these complexes display rich heterogeneity in their intensity–lifetime distributions (schlau2015single) and spectral properties (kruger2017reduced, goldsmith2010watching). It was proposed that the complex relationship between fluorescence intensity and lifetime cannot be adequately studied by traditional simple kinetic models (goldsmith2010watching). The heterogeneity and dynamics of light-harvesting complexes make them an excellent testbed for SMS clustering algorithms.
Typically, a probability distribution is used to visually reveal relevant states of an isolated SM system (schlau2015single, santoso2010characterizing, kalinin2008characterizing). These distributions can then be clustered based on the number of observed dense regions using clustering algorithms such as Gaussian mixture modeling (GMM) (mattana2022automated, rubbens2021phenogmm) or K-means clustering (wang2013lifetime), although any clustering algorithm could be used (xu2005survey, hartigan1975clustering). Each resulting cluster provides insight into the dynamics of the distinct fluorescence states and may enable inference of information such as the dwell times and switching rates between different states. There are many methods to determine these switching rates, such as the maximum-likelihood estimation of observables (schmid2016single) or inverse mean dwell time (schuler2013single), though the empirical definition of the switching rate is most universal (chung2018transition).
We have developed a general-purpose, open-source Python-based software application for SMS data analysis, called Full SMS (botha2024advanced), which performs, i.a., analysis of intensity time traces using a statistically robust change-point analysis (watkins2005detection). The software also allows statistical grouping of the resolved intensity levels through a combination of agglomerative hierarchical clustering and expectation maximization clustering (watkins2005detection), which significantly improves the fitting of intensity levels and lifetimes. The grouping, which also features a type of clustering, is technically capable of identifying physical states, although it may not be statistically plausible to use only grouping as an objective means to identify these states in a lifetime–intensity distribution. A more rigorous and versatile analysis method is required.
In this work, we present an extension of the former by introducing a robust, multistate objective clustering protocol on experimental post-processed data from the Full SMS software (botha2024advanced). Similar to Full SMS, the repository is publicly hosted on GitHub and can be accessed at https://github.com/BioPhysicsUP/Clustering-Protocol. To demonstrate the power of this clustering algorithm, we tested it on simulated data for benchmarking, followed by the following case studies: randomly oriented fluorescent dye molecules (Alexa Fluor 647) exhibiting rapid photoblinking, QD 605 data, which potentially also display multistate behavior (munoz2022memory), and two light-harvesting complexes as examples of multichromophoric systems.
Materials and Methods
Experimental Setup
The experimental setup is described in Ref. (kyeyune2019strong), with some minor modifications. Briefly, a ps pulsed supercontinuum laser (Fianium, SC400-4-PP), with a repetition rate of $40$ MHz, was used as the illumination source, where the central wavelength of excitation was selected using an acousto-optical tunable filter (Crystal Technology). The excitation beam was circularly polarized before passing through a spatial filter to isolate the $\text{TEM}_{00}$ mode, and thereafter focused into a near-diffraction-limited spot by a 1.45-NA Nikon oil-immersion objective in an epifluorescent configuration. The fluorescence photons were captured by the same objective, focused through a confocal pinhole using an appropriately selected dichroic beamsplitter, spectrally cleaned using a suitable long-pass filter, and detected by a single-photon avalanche photodiode (PD-050-CTE, Micro Photon Devices, Bolzano, Italy, IRF $\sim 128$ ps).
Sample Preparation
[…]
QD 605 ITK carboxyl-derivatized QDs (Thermo Fisher Scientific) were diluted in a solution of $10$ mM MES buffer (pH 7) with $0.2$ mM $\text{MgCl}_{2}$ and 0.050.05% (w/v) Tween-20 to a concentration of $\sim 80$ pM and sandwiched between two microscope coverslips, with the bottom one treated with poly-L-lysine (PLL). The QDs were investigated one by one using $488$ nm excitation at $140$ nW, a 605dcxt dichroic beamsplitter (Chroma Technology), and a 600LPF fluorescence filter (Edmund Optics).
[…]
Simulations
Simulated datasets were generated to benchmark the clustering algorithm, the code of which is available at https://github.com/BioPhysicsUP/SMS-Simulations. The inherent noise in SMS experiments includes the fundamental shot noise (obeying Poisson statistics), detector noise (obeying Gaussian statistics), and ambient background noise. The latter arises from stray light sources such as scattering and autofluorescence from optical components and appears as a constant background level, where each stray light source is also affected by shot noise. Background noise is especially problematic for the lowest-intensity populations, where it significantly reduces the signal-to-noise ratio. The detector noise includes dark counts and readout noise, but can be countered by taking a background measurement and subtracting that level from the intensity data. The shot noise is inherent due to the discrete nature of photons arriving at random times. This is generally the major noise contributor in experiments and cannot be subtracted. To simulate the data, random lifetime–intensity populations were generated by defining a sampling population with a user-specified mean and standard deviation (SD) for the lifetime and intensity of each state. The SD was set to $\leq 10\%$ of the corresponding mean. Dwell times were then generated on a per-particle basis: for unquenched (“on") states, they were drawn from a truncated power-law distribution (see Eq. (1)) to model heavy-tailed events arising from trap states, whereas for quenched (“off") states, a simple power-law distribution was used.
The truncated power-law function is given by
$P(t)\propto t^{-\alpha}\cdot e^{\frac{-t}{\tau_{c}}},$ (1)
where $\alpha$ is the power-law constant. The truncated power law is characterized by a high probability of short dwell-time events and a low probability of long dwell times, with $\tau_{c}$ acting as the exponential cutoff time after which the exponential decay dominates. For this particular study, we set $\alpha=1.2$ and $\tau_{c}=20$ s, and for each particle, dwell times were drawn from this distribution, where the corresponding intensities and lifetimes were sampled from the defined means and SDs, thereby simulating states with normally distributed noise. Finally, Poisson noise (representing shot noise) was added to the intensities to yield simulated two-state intensity traces.
For molecules with a distinct dipole (such as Alexa), the excitation probability depends strongly on the dipole orientation with respect to the polarization of the excitation light. In the dipole approximation, this probability is defined by
$P_{\text{Exc}}\propto|\vec{E}\cdot\vec{\mu}|^{2}\propto\sin^{2}{\theta}(\cos\phi+\epsilon\sin\phi)^{2},$ (2)
where $\theta\in[0,\pi]$ and $\phi\in[0,2\pi]$ are, respectively, the polar and azimuthal angles of the molecule’s transition dipole moment, and $\epsilon$ is the ellipticity of the excitation light’s polarization ellipse. The fluorescence quantum yield was assumed to be unaffected by the light polarization and molecule orientation.
Data analysis
The experimental data were recorded in time-tagged time-resolved (TTTR) mode, stored in HDF5 format, and analyzed using Full SMS (botha2024advanced). Data screening was the first step, whereby particles bleached (visible as an irreversible loss of fluorescence) during the raster scan were omitted, photon bursts were removed, and only the brightest particles were selected. The time traces of the remaining particles were then analyzed within a specific time region until partial or complete photodegradation was observed. The fluorescence intensity levels were then resolved and subsequently grouped using the software’s native clustering functionality, which provided the statistically most probable levels. The fluorescence lifetime of each grouped intensity level was determined using $\chi^{2}$ and Durbin-Watson goodness-of-fit (durbin1950testing) parameters.
The grouped lifetime–intensity data output of the Full SMS software was further analyzed in a raw Python environment to perform additional data clustering. This clustering algorithm (hereafter referred to as the clustering protocol) objectively determines the optimal number of identifiable populations (states) to which a single particle (object) can switch. Data were clustered using GMM, assuming that all points in the distribution were drawn from a mixture of Gaussian distributions with unknown parameters. For each GMM cluster, a mean vector and a covariance matrix were estimated, representing the cluster center and the shape/spread of the data. Data points were assigned according to the highest likelihood of belonging to a cluster. The optimal number of GMM clusters was determined using the Bayesian information criterion (BIC), a statistical measure that balances model complexity with model fit (bouman1997cluster, neath2012bayesian). The BIC value is given by
$\text{BIC}=k\cdot\text{ln}(n)-2\cdot\text{ln}(L),$ (3)
where $k$ represents the number of parameters, $n$ the number of data points, and $L$ the likelihood fitness of the model. The BIC value was determined for each number of GMM clusters, and the best fit corresponded to the lowest BIC value. Ellipses are drawn around each cluster center, each representing confidence contours for each fitted GMM component. For a two-dimensional Gaussian with mean $\vec{\mu}$ and positive semi-definite covariance matrix $\Sigma$, the contour is defined by the squared Mahalanobis distance:
$(\vec{x}-\vec{\mu})^{T}\Sigma^{-1}(\vec{x}-\vec{\mu})=\chi_{2}^{2}({p}),$ (4)
where $\vec{x}$ is a point on the lifetime–intensity plot, and $\chi^{2}_{2}(p)$ denotes the quantile of the chi-squared distribution with two degrees of freedom at probability $p$. This ellipse encloses a fraction $p$ of the probability mass around the cluster center of each GMM component. Throughout this study, the confidence level $p$ was selected empirically to minimize cluster overlap while preserving adequate information about cluster spread. Once the optimal number of populations was determined, other relevant information, such as the switching rates and switching frequencies between states, was extracted.
The switching rate from state $i$ to state $j$ was defined as
$k_{ij}=\frac{N_{ij}}{\tau_{i}},\\$ (5)
where $N_{ij}$ is the total number of switches from state $i$ to state $j$ and $\tau_{i}$ is the total dwell time in state $i$. The switching frequency indicates how frequently the particle or molecule switches between states $i$ and $j$ and was defined as
$f=\frac{N_{ij}+N_{ji}}{\tau_{i}+\tau_{j}}.$ (6)
The switching ratio, defined as
$R_{ij}=\frac{k_{ij}}{k_{ji}},$ (7)
gives the ratio of the forward to reverse switching rates and is a measure of the state preference. We used the condition $i<j$, in which case $R_{ij}<1$ indicates the sample prefers switching to the lower intensity state, $i$, and $R_{ij}>1$ indicates switching to the higher intensity state, $j$, is favored.
To make the clustering method more robust, we incorporated a modified interquartile range (IQR) outlier detection rule, which excludes data that lie sufficiently far away from the corresponding cluster center. Specifically, if the distance between the data point coordinates and the cluster center coordinates satisfies the inequality
$\text{distance}>\text{median}+k\cdot\text{IQR},$ (8)
for some constant $k$, then the data was excluded. Following the standard convention introduced by Tukey’s boxplot rule, we set $k=1.5$.
Results
Simulated Data
Simulated datasets were generated based on two-state and four-state fluorescence lifetime–intensity models to evaluate the performance of the clustering algorithm. For all sets of simulated data, the background level was chosen to be non-zero to mimic actual experiments where sources of external noise are inevitable. To benchmark the method under ideal conditions, normally distributed noise was added around the mean values of each state, constrained by their respective SDs, and these were used as the initial data from which the intensities and lifetimes were sampled. This normally distributed noise mimics the dense populations of SMS data. Specifically, for the lower intensity state, the mean and SD used for sampling of the intensity were initially set to $500\,\text{counts/s}$ (cps) and $50\,\text{cps}$, respectively, and for the higher state they were initially set to $2500\,\text{cps}$ and $150\,\text{cps}$, respectively. Under these conditions, the algorithm is expected to produce centrally clustered results that accurately reflect the underlying model structure. Figure 1 displays the results for a system of $300$ simulated particles, each having a 10-minute trace. Figure 1A shows an example of a simulated intensity trace, displaying typical photoblinking behavior, i.e., abrupt transitions between an “on" and an “off" state, where the off-state corresponds to the background noise intensity. The lifetime–intensity distribution from a set of 300 simulated traces is shown in Figure 1B, clearly revealing the two populations. The clustering protocol was applied to this set of lifetime–intensity traces, and the corresponding BIC plot is shown in Figure 1C, indicating that two clusters are identified as the statistically most favorable outcome. Figure 1D shows the optimally-clustered cluster centers in the center of each population and the 0.9 confidence ellipses using the covariance and mean for each cluster component, providing a representative illustration of the data points associated with each cluster.
Figure 1: Results of the clustering protocol performed on the two-state simulated dataset with 300 simulated particles, each with a time trace of 10 min. A. Representative example of a simulated two-state intensity trace (using $40$-ms binning for display). B. Lifetime–intensity distribution of this set of 300 simulated particles. C. Corresponding BIC score plot. D. Outcome of clustering, showing cluster centers (black crosses) and 0.9 confidence ellipses, estimated from the mean and covariance.
To incorporate the effect of the excitation light polarization and the particle orientation, we applied Eq. (2) to the different intensity states from Figure 1 for circularly polarized light ($\epsilon=1$) and a set of randomized dipole orientations. The excitation probability is dictated only by the polar angle $\theta$ in the current case (for circularly polarized light), following a $\sin^{2}$ dependence on $\theta$, i.e., the dipole orientation’s projection to the focal plane. This produces a strong broadening of the population distribution toward lower fluorescence intensities. An example intensity trace, for a particle assumed to have a fixed orientation defined by $\theta=\frac{2\pi}{3}$ and $\phi=\frac{3\pi}{4}$, yielding an excitation probability of $P_{\text{{Exc}}}=0.75$ according to Eq. (2), is shown in Figure 2A where photoblinking occurs between the off-state and the reduced on-state corresponding to a simulated maximum emission intensity of 1875 cps. The smeared lifetime–intensity distribution for a system of 300 randomly oriented particles is shown in Figure 2B, where significant broadening of the unquenched state to lower intensities, down to the background noise level, is visible. The average of all the background-subtracted intensities of the unquenched population is 830 cps, which is in agreement with the expected 1/3 statistical average of the maximum intensity of 2500 cps.
Applying the clustering protocol to this smeared dataset gives the BIC score plot in Figure 2C. The elbow points, corresponding to the points of greatest gradient increase, are at 2 and 3. These points are considered to be the first major local minima, indicating that either two or three clusters are the optimal number. The result for two clusters is shown in Figure 2D, where the cluster of the higher state is centered at its mean intensity, and that of the lower state coincides with the densely populated off-state. The ellipses around cluster centers assign the data points to the two respective clusters at a confidence level of $0.9$, providing visual insight into the cluster structure. The result for three clusters is shown in Figure 2E. The long-lifetime data are now split into two clusters, with the first cluster center (cluster 2) around the dense population at the background intensity, and the second cluster center (cluster 3) at the mean intensity of the remaining data. Visual inspection suggests that cluster 3 should rather be aligned with the population at 2500 cps. However, since the density of data at 2500 cps is significantly lower than at 500 cps, the algorithm assigns a cluster center in between the two populations, slightly closer to the higher-density population.
Notably, any other ellipticity of the excitation light’s polarization ellipse gives rise to very similar smearing of the intensity distributions down to the background level. In this representation, the ellipticity merely affects the excitation probability density, shown as a histogram in Figure 2F for equally-weighted randomly oriented dipoles interacting with three different polarization states of an incident radiation field. All three cases (linear, elliptical, and circular polarization) display a peak at an excitation probability $P_{\text{Exc}}=0$, and any non-zero choice of $\epsilon$ introduces an additional peak at $P_{\text{Exc}}=\epsilon^{2}$.
Figure 2: Influence of the excitation light polarization on the simulated intensity of a molecular dipole. A. Example intensity trace (with 40 ms binning) for circularly polarized light and a dipole orientation of $\theta=\frac{2\pi}{3}$ and $\phi=\frac{3\pi}{4}$, simulated for a time of 10 min. B. Lifetime–intensity distribution of a set of 300 fixed but randomly oriented dipoles. C. Corresponding BIC score plot. D. Cluster centers for a two-cluster model. E. Excitation probability histogram of a set of $10^{6}$ randomly oriented dipoles for circularly ($\epsilon=1$, red), elliptically ($\epsilon=0.7$, blue), and linearly ($\epsilon=0$, green) polarized light. F. Clustering results using three clusters, with cluster centers indicated by black crosses. Ellipses indicate the region of data that belongs to each cluster, at a confidence of $0.9$.
To evaluate the clustering protocol on a more complicated system, the protocol was applied to a simulated four-state system that was generated by adding two additional states, with intensities of 1200 cps and 1800 cps and corresponding SDs of 120 cps and 180 cps, respectively, to the original two-state dataset. The lifetime–intensity distribution of this dataset shows four well-defined states (Fig. S1A), which are seamlessly resolved by our clustering protocol (Fig. S1B–C). It is clear that the clustering protocol identified cluster centers corresponding exactly to the centers of the dense data regions.
Our benchmarking of the clustering protocol on simulated datasets was successful: the well-defined two- and four-state models were clustered exceptionally well with cluster centers coinciding with dense data regions. For the smeared dataset, the protocol correctly assigned cluster centers based on the underlying density of the data distribution, thus identifying the statistically most meaningful populations, despite one of the clusters appearing visually counterintuitive. In addition, the confidence ellipses not only assign the data to the most appropriate clusters but also depict the spread in the data connected to each cluster.
Alexa
We applied our clustering protocol to both the ungrouped and grouped-level Alexa data to demonstrate the effectiveness of the grouping technique described in Ref. (botha2024advanced). The Alexa dataset was chosen as a case study of rapid photoblinking, with dwell times in intensity states generally too short to be resolved visually or with a statistical change-point analysis (Fig. 3A). This was attained by deliberately not removing oxygen from the buffer solution. In addition, spin-coating of the sample introduced a random distribution of dipole orientations. We used circularly polarized excitation light to consider only the polar angle dependence of the dipole orientation, similarly to Figure 2. The resulting erratic intensity behavior and data smearing are shown in Figure 3, with the left column representing the resolved-level (ungrouped) data and the right column representing the grouped-level data. Clearly, change-point analysis of the data had difficulty resolving the actual intensity levels accurately (Fig. 3A, green trace), whereas this improved markedly after grouping the data using the built-in grouping feature of Full SMS (Fig. 3B, green trace). The lifetime–intensity distribution of the ungrouped data (Fig. 3C) was also significantly more spread out than that of the grouped data (Fig. 3D) and showed little discernible trend, making it difficult to distinguish between individual populations. In contrast, the unbiased grouping (Fig. 3D) statistically assigns similar resolved levels to a single grouped level, which effectively denoises the data and facilitates easier and more accurate clustering. Furthermore, the grouping also removed most of the short-lifetime data in the ungrouped distribution, which highlights the large fitting uncertainties when dealing with low photon budgets.
For both the ungrouped and grouped data, the BIC score plots indicate that three and five clusters correspond to the statistically most optimal values (Figs. 3E and F). The BIC score plot typically does not obey a monotonically decreasing trend and, therefore, could possess multiple local minima. Motivated by Occam’s razor, we adopt the first elbow points as the optimal solutions, which occur after the first steepest gradient descent in Figure 3E and at the first major local minimum in Figure 3F. This ensures the simplest model is selected to sufficiently describe the data without overfitting.
The resulting three-cluster analyses are shown in Figures 3G and H for the ungrouped and grouped data, respectively. The cluster centers of the ungrouped data (Fig. 3G) do not coincide with the densest populations due to the significant spread in the data; instead, the cluster centers correspond to the mean values of the data contained in the confidence ellipses. In contrast, the cluster centers of the grouped-level data correspond well to their visually expected positions (Fig. 3H). Clearly, grouping of the data leads to more streamlined and accurate clustering.
The choice of the first local minimum in the BIC analysis prevents overfitting. If the second minimum (corresponding to five clusters) is used instead (Fig. S2), state 2 in Figure 3H appears to be partitioned into multiple separate clusters, which is less likely to accurately describe the system. This inaccuracy is further highlighted by the considerably increased overlap between the confidence ellipses, creating significant ambiguity in cluster assignment.
Figure 3: Comparison of the performance of the clustering protocol on the ungrouped (left column) and grouped (right column) Alexa data with 52 particles. A and B. Example intensity traces over 10-s windows, with grey representing the 40-ms binned intensity data and green showing the ungrouped (A) or grouped (B) trace. C and D. Lifetime–intensity distributions for the corresponding datasets. E and F. BIC score plots, indicating local minima used to determine the optimal number of clusters. G and H. Clustering outcomes using three clusters, with cluster centers indicated by black crosses, with corresponding 0.85 confidence ellipses.
To eliminate extreme outliers, we implemented the modified IQR outlier detection rule (described in Eq. (8)) to identify data points that lay unusually far from their respective cluster centers. This procedure eliminated roughly $2\%$ of the data points, resulting in a retention of $\sim 98\%$, and ensured that subsequent calculations of any cluster-based statistics were not skewed by outliers, while keeping the cluster assignments unchanged. For the grouped data, clustering revealed three states with lifetimes of $0.74\pm 0.15\,\text{ns}$, $2.01\pm 0.47\,\text{ns}$, and $1.8\pm 0.1\,\text{ns}$, respectively, with an average weighted lifetime of $1.73\pm 0.55\,\text{ns}$. There are $\frac{N(N-1)}{2}$ possible reversible pairwise switching pathways for $N$ clusters. The dynamics of the three pathways between the three clusters for the Alexa data are summarized in Table 1, indicating that Alexa tends to switch preferably from state 1 to state 2 ($R_{12}>>1$), from state 3 to state 1 ($R_{13}<1$), and from state 3 to state 2 ($R_{23}<<1$).
Clusters (i,j) $k_{\text{{i}}\text{{j}}}\,(\text{s}^{-1})$ $k_{\text{{ji}}}\,(\text{s}^{-1})$ $f\,(\text{s}^{-1})$ $R_{\textit{ij}}$
(1, 2) 6.06 1.25 2.08 4.84
(1, 3) 4.29 7.99 5.59 0.54
(2, 3) 0.134 1.06 0.23 0.13
Table 1: Switching rates ($k_{ij}$), switching frequencies ($f$), and switching ratios ($R_{ij}$) between the relevant clustered states of Alexa, using Eqs. (5)-(7).
QD 605
To test the robustness of the clustering analysis to statistically small datasets, we considered a dataset comprising only 30 QDs, measured for 30 s each. QD 605 consists of a CdSe core surrounded by a ZnS shell. When the shell is sufficiently thin, these QDs, which belong to type II-IV QDs, typically exhibit bimodal blinking behavior (palstra2021python, efros2016origin). A representative intensity trace is shown in Figure 4A, illustrating the relatively fast photoblinking dynamics of this sample, although, unlike the Alexa data, the intensity levels could be well-resolved by the change-point algorithm. The lifetime–intensity distributions for the ungrouped and grouped QD data are shown in Figure 4B and C, respectively. The grouping effectively reduces the noise in the data, as evidenced by the markedly narrower lifetime–intensity distribution, which uncovers the underlying linear correlation between the two parameters. The significant denoising facilitated the subsequent data clustering. The corresponding BIC score plots show a shallow gradient from two clusters and higher for both data sets (Figs. 4 D and E), suggesting that any of two to eight clusters would be an appropriate choice, although two is the preferred number as it occurs after the first steepest gradient. The results of a two-cluster model are shown in Figures 4F and G for the ungrouped and grouped data, respectively, with ellipses indicating data assignment per cluster shown at a confidence of $0.85$.
We implemented the modified IQR outlier detection rule (Eq. (8)) on the grouped data prior to calculating the switching rates and frequencies. Approximately $4\%$ of the data points were removed, resulting in a retention of $\sim 96\%$ of the original data. This step ensured that extreme outliers were disregarded so that they would not skew the computed dynamics. The lifetimes corresponding to the two cluster centers of the grouped data are $2.2\pm 1.2\,\text{ns}$ and $9.5\pm 3.3\,\text{ns}$, with a weighted average of $7.0\pm 4.6\,\text{ns}$. The switching rates between the two states were $k_{12}=1.68\,\text{s}^{-1}$ and $k_{21}=3.04\,\text{s}^{-1}$, with a switching frequency of $f=2.16\,\text{s}^{-1}$, indicating that switching from state 2 to state 1 is more likely, i.e., state 1 is the more stable state, in agreement with its higher population density.
Figure 4: Clustering results for the statistically small QD 605 data with $\sim 30$ particles. A. Representative example intensity trace showing 40-ms binned data (grey) and resolved intensity levels (green). B and C. Lifetime–intensity distribution for the ungrouped (B) and grouped (C), respectively. D and E. Corresponding BIC score plots. F and G. Clustering both the ungrouped and grouped datasets using two clusters, with cluster centers indicated with black crosses, and ellipses (at a confidence of $0.85)$ indicating the assignment of data points to each cluster.
LHCII
The main light-harvesting complex of plants, LHCII, was chosen to test the capability of clustering a multichromophoric system that switches between different physiologically significant intensity–lifetime states and to address the debate about the actual number and identity of distinct states (schlau2015single, kruger2011fluorescence, kruger2011fluorescence2). Specifically, some studies suggest a multistate system that displays exclusively a linear correlation between intensity and lifetime, with some intermediate intensity states between the strong and weak emitting states (kruger2011fluorescence, kruger2011fluorescence2), while some other studies suggest the existence of a third, non-linear state in addition to the strongly and weakly emitting states (schlau2015single).
An example intensity trace for the grouped LHCII data is shown in Figure 5A, which illustrates photoblinking between unquenched, quenched, and intermediate intensity states. The blinking behavior is considerably less erratic than that observed for Alexa and QD 605. The ungrouped and grouped lifetime–intensity distributions are shown in Figures 5B and C, respectively, displaying a clear linear relationship that becomes more prominent when reducing the noise by means of grouping (Fig. 5C). Notably, the data shows no indication of the existence of the third, long-lifetime and low-intensity population that was reported in Ref. (schlau2015single). The BIC score plot from the clustering analysis indicates three clusters as the statistically most likely outcome for both the ungrouped (Fig. 5D) and grouped (Fig. 5E) data, and the results of clustering both datasets, using three clusters, are shown in Figures 5F and G, respectively. The cluster centers confirm the linear correlation between the lifetime and intensity for both the ungrouped and grouped data. It is clear that grouping makes the clustering more accurate, leading to tighter and better-defined clusters.
Figure 5: Clustering outcome for the LHCII data, comparing the ungrouped (left) and grouped (right) results for a system of 102 particles. A. Example intensity trace with 40-ms binned data (grey) and the resolved levels (green). B and C. Lifetime–intensity distributions for the ungrouped (B) and grouped (C) LHCII data. D and E. BIC scores plots for the ungrouped and grouped data, respectively. F and G. Corresponding clustering results using three clusters, with cluster centers indicated by black crosses, and ellipses drawn at a confidence of $0.85$.
The clustering revealed three states with lifetimes of $0.56\pm 0.18\,\text{ns}$, $2.1\pm 0.7\,\text{ns}$, and $3.56\pm 0.12\,\text{ns}$, respectively, with an average weighted lifetime of $2.7\,\pm 1.3\text{ ns}$. IQR filtering removed approximately $2\%$ of extreme outlier data points, retaining $\sim 98\%$ of the dataset. This preserved cluster assignments while ensuring no outliers contributed to the dynamics between clusters. The switching parameters, determined using Eqs. (5)-(7), are summarized in Table 2, indicating that LHCII tends to switch more regularly from state 2 to state 1 ($R_{12}<1$), from state 1 to state 3 ($R_{13}>1$), and from state 2 to state 3 ($R_{23}>1$). This points to a dynamic equilibrium between the three states that is shifted toward the unquenched state (state 3, representing the light-harvesting function of the complex) and away from the short-living, partially quenched state 2.
Clusters (i,j) $k_{\text{{i}}\text{{j}}}\,(\text{s}^{-1})$ $k_{\text{{ji}}}\,(\text{s}^{-1})$ $f\,(\text{s}^{-1})$ $R_{\textit{ij}}$
(1, 2) 0.21 0.28 0.24 0.75
(1, 3) 0.31 0.11 0.16 2.8
(2, 3) 0.39 0.11 0.17 3.54
Table 2: Switching rates ($k_{ij}$), switching frequencies ($f$), and switching ratios ($R_{ij}$) for the LHCII three-clustered grouped data, using Eqs. (5), (7).
Phycobilisome (PB)
PB from Synechocystis PCC6803 was chosen as the last case study due to its size, complexity, and biological significance. Comprising a few dozen polypeptides and a total of 396 pigments (dominguez2022structures), it can be considered the ultimate example of a large, multichromophoric system of biological importance that absorbs visible light. This photosynthetic light-harvesting complex exhibits a heterogeneous excitation energy landscape (cupellini2024energetic, kruger2019role), which results in considerable variations in its fluorescence emission, in terms of both energy and intensity (gwizdala2016controlling). Another property that puts our clustering analysis to the test is the instability of PB under in vitro conditions. Specifically, when isolated, this complex readily loses some of its subunits, invariably resulting in variations in absorption cross-section and sometimes in its fluorescence properties (gwizdala2016controlling). To address the consequent variations in fluorescence intensity, each complex’s intensity trace was normalized such that its maximum resolved intensity corresponded to the global maximum resolved intensity among all intensity traces, ensuring that each intensity trace was comparable to the next. This allowed direct comparison of intensity levels across complexes and considerably denoised the lifetime–intensity distributions. This normalization is one of the recently implemented functionalities in the Full SMS toolbox (botha2024advanced).
After the normalization and subsequent grouping, the PB data were clustered. A representative example intensity trace for the grouped PB data is shown in Figure 6A, which displays photoblinking between the relevant (normalized) intensity states. The intensity switching frequency of this sample was relatively low due to the use of a very low excitation power. The lifetime–intensity distribution of the ungrouped (Fig. 6B) and grouped (Fig. 6C) data shows significant heterogeneity, with several well-defined dense regions, which become more pronounced after grouping. The grouping again significantly denoises the data, producing better-defined data boundaries, which facilitates the clustering. Our clustering protocol was applied to both the ungrouped and grouped PB data, where for the ungrouped data, the BIC score plot (Fig. S3A) revealed five clusters as the optimal choice, whereas for the grouped data (Fig. 6D), three clusters correspond to a clear local minimum. The clustering results for the ungrouped data, using five clusters, are shown in Figure S3B, while a three-cluster model applied to the grouped data is depicted in Figure 6E. In both cases, no linear trend is evident; however, the cluster centers align well with the centers of visibly dense data regions.
For both the ungrouped and grouped data, the second local minimum corresponds to eight clusters (see Figs. S3A and 6D). Applying an eight-cluster model to the grouped data gives rise to considerable overlap between the confidence ellipses (Fig. S4), strongly suggesting overfitting. However, eight clusters are potentially a physically relevant choice for this complex (vide infra).
The IQR outlier detection rule (Eq. (8)) gave rise to $\sim 97\%$ data retention, and the three clustered states have lifetimes of $0.96\pm 0.23\,\text{ns}$, $1.5\pm 0.24\,\text{ns}$, and $1.57\pm 0.19\,\text{ns}$, respectively, with an average weighted lifetime of $1.62\,\pm 0.22$ ns. The dynamics between the clusters are summarized in Table 3, which reveals that the system tends to preferably switch from state 1 to state 2 ($R_{12}>>1$), from state 1 to 3 ($R_{13}>>1$), and from state 3 to 2 ($R_{23}<1$), indicating that state 2 is the most stable state and state 1 is the least stable.
Figure 6: Results of the clustering protocol applied to the grouped PB data with 101 measured particles. A. Example intensity trace of a PB particle, showing 40-ms binned data in grey, and the grouped intensity levels overlaid in green. B and C. Ungrouped and grouped PB lifetime–intensity distribution. D. BIC score plot for the grouped PB data. E. Corresponding clustered results for the grouped data, using three clusters, with cluster centers indicated by black crosses. Ellipses illustrate which data points are most likely associated with which cluster, at a confidence of $0.85$.
Clusters (i,j) $k_{\text{{i}}\text{{j}}}\,(\text{s}^{-1})$ $k_{\text{{ji}}}\,(\text{s}^{-1})$ $f\,(\text{s}^{-1})$ $R_{\textit{ij}}$
(1, 2) 0.075 0.008 0.014 9.38
(1, 3) 0.014 0.003 0.005 4.67
(2, 3) 0.004 0.009 0.006 0.44
Table 3: Switching rates ($k_{ij}$), switching frequencies ($f$), and switching ratios ($R_{ij}$) between the relevant clustered states of PB, using Eqs. (5)-(7)).
Discussion
Owing to the high sensitivity of SMS, any sources of heterogeneity, caused by inherent experimental noise, system composition, non-optimized experimental conditions, or data fitting uncertainties, can significantly distort the data distributions by introducing broadening. This is especially true for fluorescence intensity data, as it is generally strongly dependent on the excitation probability of the sample (kruger2011fluorescence), unlike the fluorescence lifetime. On the other hand, the resolved lifetime usually has a greater uncertainty than the resolved intensity, leading to additional broadening. Data broadening most often arises from noise and is much less likely from heterogeneity in physically meaningful states.
Skewed or unevenly distributed populations can bias clustering, causing cluster centers to be shifted away from where they might be expected through visual inspection. To address this, the grouping algorithm, implemented in the Full SMS software, helps to denoise the data and reveal any dominant underlying relationships between experimental parameters, thereby tightening distributions for interpretation. The effect is evident when comparing, for example, Figures 3C and 3D, where, in the latter case, a cleaner and more structured distribution is observed. Grouping also emphasizes the linear trend between lifetime and intensity in many SMS experiments (schlau2015single, botha2024advanced, gwizdala2016controlling, weston2002measuring, tinnefeld2001photophysical). By denoising and tightening the data, grouping makes subsequent clustering easier, more robust, and more reliable.
It is important to note that the clustering algorithm does not assess whether the clusters correspond to physically meaningful states or not; it simply partitions the data based on localized high-density regions. The clustering protocol fits the data with a mixture of Gaussian distributions, while the BIC score determines the optimal number of clusters by balancing fitness with complexity. In practice, this means that the algorithm favors simpler models unless the addition of more parameters would provide a substantial improvement to the fit. Notice that when performing clustering, the confidence ellipses representing relevant cluster regions may overlap. This is expected since GMM is a probabilistic method whose Gaussian components have indefinitely extending tails, unlike hard-clustering algorithms with strict boundaries (such as K-means). As a result, multiple clusters can contribute non-zero probabilities to the same data point, which is reflected by the overlapping ellipses. Nevertheless, each data point is assigned to the cluster for which it has the highest posterior probability, that is, the cluster to which it most likely belongs.
The BIC score curve does not necessarily display a monotonically decreasing trend, and therefore often exhibits several local minima. Although the global minimum suggests the most accurate statistical fit, it may lead to overfitting by assigning redundant cluster centers to a single broad state. Conversely, there is no objective justification for selecting any other local minimum, and the general assumption that a system can be modeled by a simple two-state model may be a crude oversimplification. To address this, we imposed the condition of considering the first relevant local minimum, following Occam’s razor. This approach describes the data objectively and sufficiently. Figure 3 illustrates this well. The BIC plot for the grouped data (Fig. 3F) displays multiple local minima. The first minimum, corresponding to three clusters, yields cluster centers that visually agree with the dense regions in the lifetime–intensity distributions (Fig. 3H). The second minimum, which corresponds to five clusters (Fig. S2), most likely produces overfitting since multiple cluster centers are now assigned to a single broad cluster (cluster 2 in Fig. 3H) with considerable overlap between the confidence ellipses. Nevertheless, it is still up to the user’s discretion to determine which of these states are physically significant. Caution should be exercised when certain assumptions are made about the number of relevant clusters, as this may lead to user bias, unless this decision is based on well-established knowledge about the system. The transitional dynamics between states can help to identify which clusters are physically relevant and which are not. For example, if a system with five identified clusters includes two that rarely share dynamics with other populations, they are likely artifacts, and the system may have been overfitted. In contrast, clusters that exhibit strong dynamic interactivity with one another are more likely to represent physically meaningful states.
Simulations
To evaluate this reasoning in a controlled manner, we benchmarked the clustering protocol using simulated datasets where the number and nature of each state are known a priori, thereby allowing direct comparison of the algorithm’s output to ground truth. The protocol was successful in clustering well-defined two- and four-state datasets (Figs. 1 and S1) with cluster centers aligning well with the centers of the densest regions of the distributions.
For the perturbed two-state dataset, the lifetime–intensity distribution (Fig. 2B) is smeared due to polarization and dipole orientation effects. Neither of the two candidate solutions in the BIC score plot assigned a data cluster near 2500 cps because its density is significantly lower than at 500 cps (cf. Fig. 2F). As a result, the algorithm—which penalizes unnecessary complexity—accepts two clusters as a sufficient choice to describe the two high-density data populations, while also suggesting three clusters as a more optimal solution, with the third cluster lying within the smeared-out region but closer to the highest-density population. This behavior of the clustering illustrates how the BIC inherently balances fit accuracy and model simplicity, accepting a simpler solution than what visual inspection may suggest, highlighting the potential bias and inaccuracy of visual approaches.
Alexa
For Alexa, the ungrouped data (Fig. 3C) appear more spread out with no obvious trend, whereas the grouped data (Fig. 3D) reveals the underlying trends due to substantial denoising. The large intensity spread in the grouped data corresponding to a lifetime of $\sim 1.8-2.0$ ns can be explained by Fig. 2, i.e., a strong broadening in the intensity distribution of the bright state is caused by the heterogeneous distribution of molecular orientations relative to the circularly polarized excitation light, causing large variations in the relative absorption cross-sections of the molecules. The absence of broadening down to the background level can be explained by user bias in the data selection: molecules that are oriented away from the focal plane interact only weakly with the incident light and, therefore, appear to be significantly dimmer than those oriented near-parallel to the focal plane. These dim signals are easily masked by the background noise and, as a result, those molecules are unlikely to be selected for further study. This skews the overall intensity distribution to brighter states, as is the case for Alexa (Fig. 3D).
The authors in Ref. (wang2013lifetime) report the lifetime of Alexa 647 to be $1.13\,\text{ns}$, which is longer than the bulk lifetime of 1 ns due to the use of glycerol. In our study, the sample was spin-coated and measured in the presence of oxygen. The average weighted single-molecule lifetime in our dataset ($1.73\pm 0.55\,\text{ns}$) deviates from that of previous studies, although this prolonged lifetime is not likely caused by the effects of different environmental conditions, but instead a result of the data grouping by the Full SMS software. Figures 3C and D indicate that the average lifetime of the quenched state (state 1) increased from $\sim 0.35$ ns for the ungrouped data (Fig. 3C) to $\sim 0.75$ ns for the grouped data (Fig. 3D). This indicates that most of the short-lifetime data in the ungrouped dataset was combined with longer-lifetime data during the grouping procedure, likely to limit the number of short events that lack statistical weight or are prone to fitting errors. This unavoidable artifact of data grouping can be limited by choosing another BIC value during the grouping analysis to ensure that a short-lifetime population is chosen as a distinct state.
Since Alexa switched more frequently from state 3 into states 1 and 2 (Table 1), states 2 and 3 may represent a single physically broadened state; however, previous studies (lin2015quantifying, yang2024electrochemically, herdly2023benchmarking)) suggested that Alexa may exhibit multiple distinct fluorescing states. These observations demonstrate the capability of our clustering protocol to successfully identify and separate physically meaningful subpopulations, confirming its robustness for complex datasets.
QD 605
We used QDs as a benchmarking sample, as it is widely used in photoblinking studies. In general, the blinking behavior of QDs is heavily influenced by factors such as their core composition (palstra2021python), shell thickness (efros2016origin), and the excitation power. These parameters influence the number and stability of accessible quasi-stable lifetime–intensity states. QD 605, used for the current study, has a CdSe core tuned to emit around 605 nm and capped with a thin ZnS shell with a thickness of <2<2 nm. Light absorption arises from excitonic transitions in the CdSe core, with multiple higher-energy transitions giving rise to significant broadening of the absorption toward the blue. The ZnS shell is transparent in this range and primarily serves to passivate any surface states and confine carriers. QDs with relatively thin shells generally follow Type-A blinking behavior (efros2016origin), which is driven by Auger recombination, i.e., recombination of an electron-hole pair that involves transfer of the energy to a different charge carrier instead of photon emission. This Type-A blinking is, therefore, associated with blinking from a lower-intensity, short-lifetime state (state 1) to a higher-intensity, longer-lifetime state (state 2), and hence the two states show a linear relation between the lifetime and intensity, in line with Figure 4C. The high density of the Auger-dominated state 1 arises because these datapoints occupy a narrow region in lifetime–intensity space, making the population appear tighter.
The results of the clustering protocol applied to the QD data indicated that two clusters were optimal (see Fig. 4E), and the weighted single-molecule average lifetime was found to be $7.3\,\pm 4.6\,\text{ns}$. Ref. (grecco2004ensemble) reports QD 605 lifetimes to be in the range of 102010-20 ns depending on excitation power and other experimental conditions, while the bulk (solution-phase) lifetime was measured as $6.3\pm 1.2\,\text{ns}$ in Ref. (bhuckory2016evaluating). Our experimental conditions were similar to those in the latter study, with QDs dispersed, no oxygen scavengers or additional additives such as glycerol used. Therefore, our measured average lifetime is in good agreement with the bulk value. The relatively high uncertainty in our lifetime arises primarily from the statistically small size of the dataset and the heterogeneity of the SM events. To fully establish the exact number of relevant states for this QD sample, one would have to systematically compare the switching dynamics and blinking behavior for different choices of the cluster number. Nevertheless, the two-cluster fitting is sufficient based on the switching dynamics (switching rates and frequencies) between these two clusters, which indicates a strong reversible blinking mechanism.
LHCII
LHCII is often modeled as a two-state system (gruber2018isolated, kruger2011fluorescence, kruger2011fluorescence2), though intermediate states have also been identified (kruger2011fluorescence, kruger2011fluorescence2). The LHCII sample analyzed in this study displayed a linear correlation between the intensity and lifetime, and the clustering protocol revealed a strongly emitting unquenched state, a quenched state, as well as an intermediate state (Fig. 5), consistent with previous studies (kruger2011fluorescence, kruger2011fluorescence2). The environmental conditions in our study closely matched those of Ref. (akhtar2015pigment), which reports ensemble measurements in solution without the use of oxygen scavengers, and found a bulk lifetime of 3.6ns3.6\ $\text${ns}, in good agreement with our measured long-lifetime component. In bulk studies, the average intensity per complex is orders of magnitude less than in SMS studies. Since blinking is strongly light-dependent (kruger2011fluorescence), it will contribute negligibly to the bulk lifetime, explaining why the bulk lifetime is expected to match the long-lifetime component in our study. The intermediate state (state $2$) is identified as the least stable of the three states (Table 2). If this sample were to be treated as a two-state model out of the box, it would distribute this intermediate data into both states and skew the data, indicating a perfect example of user bias. Instead, the clustering protocol, which helps with considerable noise reduction, reveals the underlying structure more accurately. This allows a user to potentially eliminate unwanted or weakly contributing clusters and analyze only the relevant clusters as if they were part of a robust, error-resistant measurement system.
PB
PB is a massive, multi-hundred-chromophoric system that is highly sensitive to excitation power and experimental conditions (gwizdala2016controlling). Since there was a large variation in the fluorescence intensities, the sample data had to be normalized to a global level prior to clustering. The excitation power was relatively low ($\sim 1.7$ nW), which led to a low fluorescence intensity switching rate (assefa2024apce), as evidenced from Table 3. The BIC analysis provides sufficient evidence that multiple clusters can be identified, with three clusters being the simplest model to describe the system. This, however, should be regarded as a minimal representation, since the lifetime–intensity distribution reveals much more complexity than only three populations, suggesting that the dynamics are far richer. In fact, eight subpopulations, the next candidate solution, are consistent with the compartmental model of van Stokkum and coworkers, where eight functional compartments were used to describe the excitation energy transfer dynamics in the core units of PB (see Fig. 9 in Ref. (van2018functional)). Importantly, their model was fully consistent with their time-resolved emission spectroscopy data. Information about higher-energy compartments can be assessed in particular for strong energy traps, in which case the emission is trap-limited and the fluorescence spectra are strongly blue-shifted (gwizdala2016controlling). Such trap-limited, pre-equilibration emission explains the clusters that show a non-linear relationship between the lifetime and intensity in Figure 6. However, this non-linearity is also partly attributed to the size heterogeneity that was not fully accounted for by our data normalization procedure. It, therefore, appears that eight clusters may be physically relevant for PB.
The longest measured lifetime component of the unquenched states is that of cluster 3, with an average weighted lifetime of $1.57\pm 0.19\,\text{ns}$, which matches well with the bulk lifetime of $\sim 1.6$ ns reported in Refs. (gwizdala2016controlling, squires2019single), where experimental conditions were similar. The expected similarity between the bulk lifetime and the long-lifetime component in our study is explained as for LHCII above, considering that the fluorescence intensity dynamics of PB are also strongly dependent on the excitation intensity (gwizdala2016controlling). The densest data of the long-lifetime component cluster, enclosed by the confidence ellipse, has a slightly longer lifetime ($1.57\pm 0.19\,\text{ns}$) than that of the cluster center ($\sim 1.5$ ns), and therefore marginally biases the average weighted lifetime upwards.
Conclusions
While highly sensitive, SMS is prone to significant noise, which can easily introduce fitting errors, bias in the parameter estimates, and artificially inflated variability. This reduces the ability to identify true subpopulations, leading to higher uncertainty and weaker correlations. It is, therefore, crucial to perform careful and reliable noise reduction to preserve meaningful structure while minimizing the spurious effects of noise, without introducing new biases.
In this work, we highlighted the comparative advantages of using grouped levels over ungrouped resolved levels in data analysis, particularly for denoising and subpopulation identification. We also introduced an objective GMM clustering protocol based on a BIC scoring system to determine the optimal number of dense clustered regions in 2D SMS data. We specifically tested this clustering protocol on SM lifetime vs. intensity data, whereby we first benchmarked the capability of the protocol on simulated two- and four-state data, taking large intensity heterogeneity into account that may result from a case such as varying excitation probabilities of randomly oriented molecular dipoles. Thereafter, we tested the clustering protocol on Alexa 647 dye, QD 605, LHCII, and PB. We identified three data clusters for Alexa, LHCII, and PB, and two clusters for QD. We also deduced that the clustering protocol is useful in the identification (and thereby the removal) of clusters corresponding to misfits and nonoptimal sample preparation, and the assignment of minor clusters that are often obscured by noise. The clustering protocol typically reveals a non-monotonically decreasing plot for the BIC score per cluster, and therefore, several minima are expected. It is therefore recommended that a separate study be done to determine the effect of different identified optimal clusters on the switching dynamics.
We have considered only pairwise switching dynamics between statistically significant subpopulations, for which the number of switching pathways scales with $\frac{N(N-1)}{2}$. An alternative analysis could involve switches between any one state and all the other N1N-1 states together, which would decrease the number of switching rates to $2N$.
Although this study focused only on fluorescence lifetime–intensity correlations, the general nature of the clustering protocol makes it applicable to any two-dimensional parameter correlations. In addition, the protocol can readily be extended to higher dimensions.
Overall, it appears that grouping the resolved levels is necessary when the data is significantly broadened, and clustering further is even more necessary to identify physically relevant subpopulations. It is clear that the clustering protocol is capable of detecting clusters for erratic data (such as the Alexa data), statistically small data sets (such as with the QD data), as well as multistate data (LHCII and PB).
Supporting Material
The Supplementary Information contains additional supportive Figures.
Acknowledgments
We wish to acknowledge Michal Gwizdala for the PB and LHCII isolations and the fruitful discussions. MACL was supported by the National Research Foundation (NRF), South Africa (PMDS22070633262). JLB was supported by the Vrije Universiteit Amsterdam-NRF Desmond Tutu programme. TPJK acknowledges funding from the NRF (grant nos. 137973 and 0403211945) and the Rental Pool Programme of the Council for Scientific and Industrial Research’s Photonics Centre, South Africa.
Author Contributions
All authors conceived the study. JLB and GTA performed the experiments, captured the data, and performed pre-clustering analyses. MACL developed the clustering software, performed the simulations, and analyzed the data. TPJK acquired funding, supervised research, and administered the project. MACL and TPJK wrote the manuscript. All authors read and approved the final manuscript.
Declaration of interests
The authors declare no competing interests.
This Supplemental Information contains additional Figures as referenced in the main text.
Supplementary Information
Simulated data additonal information
Figure S1: Outcome of the clustering protocol applied to the well-defined four-state data. A. Lifetime–intensity distribution of the data. B. Corresponding BIC score plot. C. Clustered four-state data with cluster centers indicated by black crosses. Ellipses are drawn around cluster centers, with a confidence of $0.9$.
Alexa Fluor additional information
Figure S2: Applying clustering protocol to the Alexa grouped data, using 5 clusters, to show the effect of overfitting. Ellipses are shown, with a confidence of $0.85$.
PB additional information
Figure S3: Outcome of applying the clustering protocol to the ungrouped data. A. BIC score plot for the ungrouped PB data. B. Corresponding clustering result using five clusters, with cluster centers indicated using black crosses. Representative ellipses are drawn to show which clusters correspond to which data, at a confidence of $0.85$.
Figure S4: Result of clustering the grouped PB data using eight clusters (cluster centers are shown as black crosses). Ellipses indicate the region of data points considered for each cluster, at a confidence of $0.85.$